OpenStudy (anonymous):

Math Analysis: Use Demoivre's theorem to simplify the expression. Write the answer in a + bi form (4cis135)^4

OpenStudy (anonymous):

(4cis135)^4=4^4(cos(135*4))

OpenStudy (anonymous):

is the answer -256?

OpenStudy (anonymous):

256(cos180)=-256

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

what about sine b/c it's "cis"

OpenStudy (anonymous):

i tought cis was a typo

OpenStudy (anonymous):

cis means cos(theta)+i*sin(theta)

OpenStudy (anonymous):

yea so the answer would just be -256 + 0i?

OpenStudy (anonymous):

oh.. then 256(cos180 +i sin180)=-256

OpenStudy (anonymous):

ya, it's pure real number

OpenStudy (anonymous):

Where is the 180 coming from?

OpenStudy (anonymous):

cos(135)=-root(2)/2

OpenStudy (anonymous):

idk. all i know is that the answer is -256

OpenStudy (anonymous):

no, from my teacher's work it's -256

OpenStudy (anonymous):

No, I meant cos135=+root(2)/2

OpenStudy (anonymous):

it wants the answer in a + bi form so i guess the 180 comes from multiplying the 135 by 4? o-o

OpenStudy (anonymous):

135)^4 which is really 135 x 4 is 180

OpenStudy (anonymous):

Hm...I've never seen that conversion to multiplication before

OpenStudy (anonymous):

135*4=540 540-360=180

OpenStudy (anonymous):

i know.. :\ idk but the right answer is -256 from my teacher's handouts

OpenStudy (anonymous):

oh yeah, sorry lol

OpenStudy (anonymous):

cos is a periodic function. So doesn't matter is angle is 180, 540, 900..... all same values of cos

OpenStudy (anonymous):

Right right, but wasn't the original problem \[(4cis135)^4\]

OpenStudy (anonymous):

So that means you would need to evaluate cis(135) and then multiply by 4

OpenStudy (anonymous):

don't confuse him. The proceedure is right. The angle is multiplied, not the cos of the angle!!

OpenStudy (anonymous):

4cis(135) doesn't mean that the angle is multiplied...

OpenStudy (anonymous):

this is a polar expretion of a comlex number

OpenStudy (anonymous):

5cos(0)=5 cos(5*0)=1 It's not the same thing.

OpenStudy (anonymous):

(4cis135)^4 is the problem

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Demoivre's theorem look in google

OpenStudy (anonymous):

Ah, quite right. My bad.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Apologies.

OpenStudy (anonymous):

Write the complex number in trigonometric form, using degree measure for the argument. 23. -10 - 24i

OpenStudy (anonymous):

lol might as well post up my next question here o-o

OpenStudy (anonymous):

you should post a new question in new window

OpenStudy (anonymous):

gg k :)